# DeepOHeat-v1: Efficient Operator Learning for Fast and Trustworthy Thermal Simulation and Optimization in 3D-IC Design

### v1

@@ Abstract
Thermal analysis is crucial in three-dimensional integrated circuit (3D-IC) design due to increased power density and complex heat dissipation paths. Although operator learning frameworks such as DeepOHeat [1] have demonstrated promising preliminary results in accelerating thermal simulation, they face critical limitations in prediction capability for multi-scale thermal patterns, training efficiency, and trustworthiness of results during design optimization. This paper presents DeepOHeat-v1, an enhanced physics-informed operator learning framework that addresses these challenges through three key innovations. First, we integrate Kolmogorov-Arnold Networks with learnable activation functions as trunk networks, enabling an adaptive representation of multi-scale thermal patterns. This approach achieves a 1.25×1.25$\times$1.25 × and 6.29×6.29$\times$6.29 × reduction in error in two representative test cases. Second, we introduce a separable training method that decomposes the basis function along the coordinate axes, achieving 62×62$\times$62 × training speedup and 31×31$\times$31 × GPU memory reduction in our baseline case, and enabling thermal analysis at resolutions previously infeasible due to GPU memory constraints. Third, we propose a confidence score to evaluate the trustworthiness of the predicted results, and further develop a hybrid optimization workflow that combines operator learning with finite difference (FD) using Generalized Minimal Residual (GMRES) method for incremental solution refinement, enabling efficient and trustworthy thermal optimization. Experimental results demonstrate that DeepOHeat-v1 achieves accuracy comparable to optimization using high-fidelity finite difference solvers, while speeding up the entire optimization process by 70.6×70.6$\times$70.6 × in our test cases, effectively minimizing the peak temperature through optimal placement of heat-generating components.
Index Terms:
3D IC, thermal simulation, operator learning, kolmogorov–arnold networks.
@@ I Introduction
Three-dimensional integrated circuit (3D-IC) is a promising solution to address the scaling challenges of semiconductor chips[2]. Although vertical stacking of multiple active silicon layers enables a higher integration density and improved system performance [3], it also introduces critical thermal management challenges [4, 5, 6]. Heat generated in one layer must pass through multiple layers before reaching the heat sink, creating thermal bottlenecks and potential reliability issues [7]. Addressing these thermal challenges requires extensive analysis and optimization during the design phase. This process involves solving partial differential equations for numerous design configurations, making it computationally intensive. Traditional numerical methods, such as finite element method and finite difference method, while accurate, are too time-consuming for iterative design optimization [6, 8, 9]. This has motivated the development of various surrogate models, including model order reduction [10, 11] and data-driven approaches [12, 13, 14].
@@
  • Accuracy improvement through Kolmogorov-Arnold networks (KAN). Unlike MLPs with fixed activation functions, KANs employ learnable univariate functions[22, 23], allowing a more adaptive representation of multiscale thermal patterns. We express KAN as DeepOHeat trunk networks, achieving 1.25×1.25$\times$1.25 × and 6.29×6.29$\times$6.29 × error reductions in two representative test cases.
  • Efficiency improvement via separable training. We employ recently developed separable training method [24] to decompose the basis functions of DeepOHeat along coordinate axes. Using forward-mode automatic differentiation [25], this approach achieves 62×62$\times$62 × training speedup and 31×31$\times$31 × memory reduction in our baseline case, making high-resolution thermal analysis computationally feasible.
  • Trustworthy thermal optimization. Lack of trustworthiness is a fundamental challenge that prevents the application of almost all neural network-based surrogates in design optimization. We propose a confidence estimator to evaluate whether a DeepOHeat prediction is acceptable or not at every step of an optimization loop. Once a prediction is not acceptable (for some corner design cases), we use it as an initial guess and call GMRES (the Generalized Minimal Residual Method [26]) for only a small number of iterations to get a high-fidelity finite-difference (FD) solution. This hybrid framework achieves an excellent trade-off between computational efficiency and solution accuracy. We apply this optimization framework to optimize the floorplan of 3D-IC for improved thermal characteristics and obtain an accuracy comparable to that of an FD-based optimizer while achieving 70.6×70.6$\times$70.6 × speedup.
@@ II-A Thermal Simulation in 3D-ICs
Under steady-state conditions ($\frac{\partial T}{\partial t}=0$), Equation (1) reduces to:
@@
    $-k(y_{3})\frac{\partial T}{\partial y_{i}}=q_{n},$ (3)
@@
    $-k(y_{3})\cdot\frac{\partial T}{\partial y_{i}}=h(T-T_{\rm amb}),$ (5)
@@ III Kolmogorov–Arnold Networks with Multi-Frequency Basis Functions
A major challenge in learning thermal fields is to capture both smooth varying regions and sharply localized hotspots. Multilayer perceptrons (MLPs) often struggle with such multi-scale patterns due to spectral bias [20], preferring low-frequency solutions over high-frequency ones [31]. To address this, DeepOHeat-v1 uses a Kolmogorov–Arnold Network (KAN) [22] as a trunk network. Kolmogorov’s representation theorem [33] states that any continuous multivariate function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ can be represented as:
@@ III-B Theoretical Understanding of the Training Dynamics
where $a_{k}$’s are trainable parameters and $C_{k}(y)$’s are the Chebyshev polynomials defined in Eq. (14). Our network approximates the first $K\!+\!1$ components of $f^{*}$ through gradient descent. Understanding this multi-frequency learning process provides insights into the behavior of deeper KAN architectures.
@@
Under gradient flow (see Appendix -A), the output evolves as:
@@
where $t$ is the training time and $\eta>0$ is the learning rate. The error $e(y,t)=f(y,t)-f^{*}(y)$ decays spectrally (see Appendix -A):
@@ V Trustworthy Design Optimization
In this section, we demonstrate how DeepOHeat-v1 can be integrated into a trustworthy design optimization framework for 3D-IC thermal optimization.
@@ V-C A Trustworthy Hybrid Optimization Framework
In this work, we use simulated annealing [36] as the outer-loop optimizer (see Algorithm 1 in Appendix -B). However, our adaptive solution method can be integrated with any other outer-loop optimizer, including but not limited to gradient descent, reinforcement learning, and Bayesian optimization.
@@ VI Experiments
We evaluate our enhanced framework through comprehensive experiments, systematically demonstrating how each proposed improvement addresses the limitations of DeepOHeat. First, we assess the framework on a 2D power map thermal simulation benchmark to validate both prediction accuracy and computational efficiency improvements. Second, we validate our trustworthy optimization workflow in a floorplan design scenario, where reliable temperature prediction is crucial for finding optimal component placements. Throughout our experiments, we compare following models:
@@ VI-A1 Problem Formulation and Implementation
All models share the same branch network architecture (9 layers, 256 neurons per layer) processing the 441-dimensional flattened power map input. For trunk networks, the baseline DeepOHeat employs a single 6-layer MLP (128 nodes/layer) with Fourier feature mapping ($\mu=0$, $\sigma=2\pi$) and swish activation. DeepOHeat + KAN maintains the single trunk network architecture but replaces the MLP with a 4-layer order-3 Chebyshev KAN (64 nodes/layer) to keep the parameter budget comparable. DeepOHeat-v1 implements our separable training approach with three independent trunk networks (one per spatial dimension), each having the same KAN architecture as DeepOHeat + KAN. The physics-informed loss is evaluated on the full computational mesh. To generate diverse training data, we sample 50 power maps per iteration from a two-dimensional standard Gaussian random field (GRF) with length scale parameter of 0.3. All implementations are built with JAX for efficient computation. Models are trained for 10,000 iterations using ADAM optimizer (learning rate 1e-3, decayed by 0.9$\times$ every 500 iterations) on an NVIDIA RTX 3090 GPU.
@@ VI-A2 Results Analysis
Table I presents performance comparisons with $N_{f}=50$ sampled functions per iteration. The reported mean absolute percentage error (MAPE) values are calculated and averaged based on ten representative power maps from the original DeepOHeat paper. DeepOHeat + KAN demonstrates superior prediction accuracy, reducing MAPE from 0.060% to 0.048% compared to the baseline. This 1.25×1.25$\times$1.25 × reduction in error validates that KAN’s learnable activation functions better capture multi-scale thermal patterns. The computational cost remains comparable to the baseline (0.51h vs 0.53h, 17.32GB vs 17.44GB), indicating the improved accuracy comes without additional overhead.
DeepOHeat-v1 maintains the enhanced prediction capability (0.049% MAPE) while drastically reducing computational requirements through separable training. Training time decreases by 62×62$\times$62 × (from 0.51h to 30.24s) and memory usage by 31×31$\times$31 × (from 17.32GB to 0.56GB), demonstrating effective mitigation of the computational bottleneck.
@@ VI-B2 Surrogate Model Training and Evaluation
The high-resolution mesh ($101\times 101\times 56$) presents a significant computational challenge, requiring more than 570K spatial points during the physics-informed training. This scale makes the baseline DeepOHeat computationally infeasible due to memory limitations. Only models incorporating separable training—DeepOHeat + ST and DeepOHeat-v1—can be effectively trained on an NVIDIA RTX 3090 GPU.
We maintain the same network architectures as in Section VI-A, with the only difference being the branch network input dimension increased to 10,201 to handle the $101\times 101$ power map in the middle layer. During training, we generate random floorplan configurations by stochastically placing the ten components, allowing overlap for efficient sampling. Each model is trained for 100,000 iterations using ADAM optimizer (learning rate 1e-3, decayed by 0.9$\times$ every 1,000 iterations), with 50 floorplans sampled per iteration. The physics-informed loss enforces thermal governing equations and boundary conditions across the entire computational domain.
For performance evaluation, we generate 100 random test floorplans following the same sampling procedure as the training data. The prediction accuracy is measured against reference solutions computed using Generalized Minimal Residual Method (GMRES) with algebraic multigrid preconditioning. Table II presents a detailed performance comparison. DeepOHeat-v1 achieves dramatically improved accuracy, reducing MAPE from 0.22% to 0.035%—a 6.29$\times$ reduction in error—while maintaining similar computational efficiency (approximately 0.5 hours of training time and 12GB memory usage). This significant improvement highlights KAN’s superior capability in capturing complex 3D thermal patterns, while separable training enables scalability to high resolutions.
@@ VI-C2 Overall Performance of Thermal Optimization for Floorplan
We evaluate our hybrid framework in a thermal optimization task in floorplanning. The optimization objective is to minimize the peak temperature of the chip model by determining optimal component placements while avoiding overlaps through adding a soft penalty term to the objective function.We implement simulated annealing using Optuna [38] with 1,000 optimization steps, an initial SA temperature of 0.5, a cooling rate of 0.997, and a base neighborhood size of 25 (as detailed in Algorithm 1 in Appendix -B). For the hybrid framework’s residual threshold, we set $\alpha$ to 9.
@@
Table III compares the three thermal evaluation approaches. The results demonstrate the effectiveness of our hybrid framework in balancing accuracy and efficiency. GMRES with algebraic multigrid preconditioning serves as our golden reference, providing the most accurate solution but at the highest computational cost. Pure operator learning achieves significant speedup (0.05h vs. 4.94h), but does not offer a guarantee of trustworthiness during the optimization process. Our hybrid approach addresses this concern by monitoring the quality of the prediction through residual verification, triggering non-preconditioned GMRES refinement in 299 out of 1,000 optimization steps. By ensuring reliable temperature evaluation throughout the optimization, the hybrid approach not only provides solution guarantees but also achieves better design quality (327.31$(K)$ vs. 328.50$(K)$ maximum temperature) compared to pure operator learning, while closely matching the reference solution (327.18$(K)$). This improvement comes with minimal computational overhead (0.02h additional optimization time), demonstrating the value of selective refinement for trustworthy design optimization.
The final optimized floorplans in Fig. 14 show that our hybrid approach produces design solutions of similar quality to non-machine learning method while reducing computating timet by 70×70$\times$70 ×.
@@ VII Conclusion
This work has presented DeepOHeat-v1 to significantly enhance the original DeepOHeat framework through three key innovations. By replacing MLP trunk networks with Kolmogorov-Arnold Networks, we have enabled adaptive learning of multi-scale thermal features, leading to substantial improvements in prediction accuracy for both surface and volumetric power distributions. The employment of separable training has overcome the computational bottleneck in physics-informed training, making high-resolution thermal analysis feasible on standard GPU hardware. Most importantly, our confidence estimator and hybrid optimization framework transform operator learning from a fast but potentially unreliable predictor into a trustworthy design optimization tool for thermal optimization in 3D-IC design. The selective refinement strategy effectively balances efficiency and reliability, achieving accuracy comparable to an optimization engine using gold-reference mesh-based PDE solvers while obtaining significant speedup. This capability is particularly valuable for practical 3D-IC design, where reliable thermal evaluation is crucial to finding optimal designs in design space exploration and signal/power integrity sign-off.
These advances open new possibilities for efficient thermal optimization in increasingly complex 3D-IC and chiplet design. In future work, we plan to extend our framework to address more practical thermal-aware floorplan optimization problems, including considerations for power delivery networks, timing constraints, and multi-objective optimization scenarios that balance thermal performance with other critical design metrics.

### v2

@@ Abstract
Thermal analysis is crucial in three-dimensional integrated circuit (3D-IC) design due to increased power density and complex heat dissipation paths. Although operator learning frameworks such as DeepOHeat [1] have demonstrated promising preliminary results in accelerating thermal simulation, they face critical limitations in prediction capability for multi-scale thermal patterns, training efficiency, and trustworthiness of results during design optimization. This paper presents DeepOHeat-v1, an enhanced physics-informed operator learning framework that addresses these challenges through three key innovations. First, we integrate Kolmogorov-Arnold Networks with learnable activation functions as trunk networks, enabling an adaptive representation of multi-scale thermal patterns. This approach achieves a $1.25\times$ and $6.29\times$ reduction in error in two representative test cases. Second, we introduce a separable training method that decomposes the basis function along the coordinate axes, achieving $62\times$ training speedup and $31\times$ GPU memory reduction in our baseline case, and enabling thermal analysis at resolutions previously infeasible due to GPU memory constraints. Third, we propose a confidence score to evaluate the trustworthiness of the predicted results, and further develop a hybrid optimization workflow that combines operator learning with finite difference (FD) using Generalized Minimal Residual (GMRES) method for incremental solution refinement, enabling efficient and trustworthy thermal optimization. Experimental results demonstrate that DeepOHeat-v1 achieves accuracy comparable to optimization using high-fidelity finite difference solvers, while speeding up the entire optimization process by $70.6\times$ in our test cases, effectively minimizing the peak temperature through optimal placement of heat-generating components. Open source code is available at https://github.com/xlyu0127/DeepOHeat-v1.
@@ I Introduction
Three-dimensional integrated circuit (3D-IC) is a promising solution to address the scaling challenges of semiconductor chips[2]. Although vertical stacking of multiple active silicon layers enables a higher integration density and improved system performance [3], it also introduces critical thermal management challenges [4, 5, 6]. Heat generated in one layer must pass through multiple layers before reaching the heat sink, creating thermal bottlenecks and potential reliability issues [7]. Addressing these thermal challenges requires extensive analysis and optimization during the design phase. This process involves solving partial differential equations for numerous design configurations, making it computationally intensive. Traditional numerical methods, such as finite element method and finite difference method, while accurate, are too time-consuming for iterative design optimization [6, 8, 9]. This has motivated the development of various surrogate models, including model order reduction [10, 11] and data-driven approaches [12, 13, 14]. Recent advances in multiscale thermal modeling, such as the ROM-based approach by [15], have demonstrated the importance of capturing disparate length scales in 3D-IC thermal analysis.
@@
  • Accuracy improvement through Kolmogorov-Arnold networks (KAN). Unlike MLPs with fixed activation functions, KANs employ learnable univariate functions[23, 24], allowing a more adaptive representation of multiscale thermal patterns. We express KAN as DeepOHeat trunk networks, achieving $1.25\times$ and $6.29\times$ error reductions in two representative test cases.
  • Efficiency improvement via separable training. We employ recently developed separable training method [25] to decompose the basis functions of DeepOHeat along coordinate axes. Using forward-mode automatic differentiation [26], this approach achieves $62\times$ training speedup and $31\times$ memory reduction in our baseline case, making high-resolution thermal analysis computationally feasible.
  • Trustworthy thermal optimization. Lack of trustworthiness is a fundamental challenge that prevents the application of almost all neural network-based surrogates in design optimization. We propose a confidence estimator to evaluate whether a DeepOHeat prediction is acceptable or not at every step of an optimization loop. Once a prediction is not acceptable (for some corner design cases), we use it as an initial guess and call GMRES (the Generalized Minimal Residual Method [27]) for only a small number of iterations to get a high-fidelity finite-difference (FD) solution. This hybrid framework achieves an excellent trade-off between computational efficiency and solution accuracy. We apply this optimization framework to optimize thermal characteristics in 3D chip designs and obtain an accuracy comparable to that of an FD-based optimizer while achieving $70.6\times$ speedup.
@@ II-A Thermal Simulation in 3D-ICs
In this work, we focus on steady-state thermal analysis. Under steady-state conditions ($\frac{\partial T}{\partial t}=0$), Equation (1) reduces to:
@@
    $k(y_{3})\frac{\partial T}{\partial y_{i}}=q_{n},$ (3)
@@
    $k(y_{3})\cdot\frac{\partial T}{\partial y_{i}}=h(T-T_{\rm amb}),$ (5)
@@ III Kolmogorov–Arnold Networks with Multi-Frequency Basis Functions
A major challenge in learning thermal fields is to capture both smooth varying regions and sharply localized hotspots. Multilayer perceptrons (MLPs) often struggle with such multi-scale patterns due to spectral bias [21], preferring low-frequency solutions over high-frequency ones [32]. In this context, ’multi-scale’ refers to thermal fields that simultaneously contain low-frequency spatial components (smooth, gradually-varying regions) and high-frequency components (sharp temperature gradients and localized hotspots). The challenge lies not in having features of vastly different sizes, but in accurately representing both the smooth and sharp frequency characteristics that coexist in thermal solutions, particularly around heat sources and material interfaces where sharp gradients naturally occur even with similar-magnitude heating. To address this, DeepOHeat-v1 uses a Kolmogorov–Arnold Network (KAN) [23] as a trunk network. Kolmogorov’s representation theorem [34] states that any continuous multivariate function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ can be represented as:
@@ III-B Theoretical Understanding of the Training Dynamics
where $a_{k}$’s are trainable parameters and $C_{k}(y)$’s are the Chebyshev polynomials defined in Eq. (14). Our network approximates the first K+1K\!+\!1 components of $f^{*}$ through gradient descent. Understanding this multi-frequency learning process provides insights into the behavior of deeper KAN architectures.
@@
Under gradient flow (see Appendix -B), the output evolves as:
@@
where $t$ is the training time and $\eta>0$ is the learning rate. The error $e(y,t)=f(y,t)-f^{*}(y)$ decays spectrally (see Appendix -B):
@@ V Trustworthy Design Optimization
In this section, we demonstrate how DeepOHeat-v1 can be integrated into a trustworthy design optimization framework for thermal optimization. By “trustworthy”, we mean predictions with quantifiable confidence and mechanisms for verification and refinement when accuracy is insufficient.
@@ V-C A Trustworthy Hybrid Optimization Framework
In this work, we use simulated annealing [37] as the outer-loop optimizer (see Algorithm 1 in Appendix -C). However, our adaptive solution method can be integrated with any other outer-loop optimizer, including but not limited to gradient descent, reinforcement learning, and Bayesian optimization.
@@ VI Experiments
We evaluate our enhanced framework through comprehensive experiments, systematically demonstrating how each proposed improvement addresses the limitations of DeepOHeat. Our experiments use $1\text{mm}\times 1\text{mm}$ chip dimensions in the $xy$-direction, following the original DeepOHeat paper for fair comparison, while our KAN and separable training improvements are architecture-independent and provide greater benefits at larger scales. First, we assess the framework on a 2D power map thermal simulation benchmark to validate both prediction accuracy and computational efficiency improvements. Second, we validate our trustworthy optimization workflow in a floorplan design scenario, where reliable temperature prediction is crucial for finding optimal component placements. Throughout our experiments, we compare following models:
@@ VI-A1 Problem Formulation and Implementation
[…]
@@ VI-A2 Results Analysis
Table I presents performance comparisons with $N_{f}=50$ sampled functions per iteration. The reported mean absolute percentage error (MAPE) values are calculated and averaged based on ten representative power maps from the original DeepOHeat paper. Appendix -A provides visualizations of all 10 test cases. Importantly, these evaluation power maps consist of structured heat blocks that differ significantly from the Gaussian Random Field (GRF) patterns used during training. DeepOHeat + KAN demonstrates superior prediction accuracy, reducing MAPE from 0.060% to 0.048% compared to the baseline. This $1.25\times$ reduction in error validates that KAN’s learnable activation functions better capture multi-scale thermal patterns. The computational cost remains comparable to the baseline (0.51h vs 0.53h, 17.32GB vs 17.44GB), indicating the improved accuracy comes without additional overhead.
DeepOHeat-v1 maintains the enhanced prediction capability (0.049% MAPE) while drastically reducing computational requirements through separable training. Training time decreases by $62\times$ (from 0.51h to 30.24s) and memory usage by $31\times$ (from 17.32GB to 0.56GB), demonstrating effective mitigation of the computational bottleneck.
@@ VI-B2 Surrogate Model Training and Evaluation
The high-resolution mesh ($101\times 101\times 56$), representing a spatial resolution of approximately 10 µm per grid point, presents a significant computational challenge, requiring more than 570K spatial points during the physics-informed training. During training, we minimize the physics loss across this entire fine mesh, enforcing thermal governing equations and boundary conditions at high spatial resolution. This high-resolution physics-informed training produces more accurate surrogate models compared to training on coarser grids, and was computationally infeasible with the original DeepOHeat due to memory limitations. Only models incorporating separable training—DeepOHeat + ST and DeepOHeat-v1—can be effectively trained on an NVIDIA RTX 3090 GPU.
[…]
@@ VI-C2 Overall Performance of Thermal Optimization for Floorplan
We evaluate our hybrid framework in a thermal optimization task in floorplanning. The optimization objective is to minimize the peak temperature of the chip model by determining optimal component placements while avoiding overlaps through adding a soft penalty term to the objective function.We implement simulated annealing using Optuna [39] with 1,000 optimization steps, an initial SA temperature of 0.5, a cooling rate of 0.997, and a base neighborhood size of 25 (as detailed in Algorithm 1 in Appendix -C). For the hybrid framework’s residual threshold, we set $\alpha$ to 9.
@@
Table III compares the three thermal evaluation approaches. The results demonstrate the effectiveness of our hybrid framework in balancing accuracy and efficiency. GMRES with algebraic multigrid preconditioning serves as our reference, providing the most accurate solution but at the highest computational cost. Pure operator learning achieves significant speedup (0.05h vs. 4.94h), but does not offer a guarantee of trustworthiness during the optimization process. Our hybrid approach addresses this concern by monitoring the quality of the prediction through residual verification, triggering non-preconditioned GMRES refinement in 299 out of 1,000 optimization steps. By ensuring reliable temperature evaluation throughout the optimization, the hybrid approach not only provides solution guarantees but also achieves better design quality (327.31(K)(K) vs. 328.50(K)(K) maximum temperature) compared to pure operator learning, while closely matching the reference solution (327.18(K)(K)). This improvement comes with minimal computational overhead (0.02h additional optimization time), demonstrating the value of selective refinement for trustworthy design optimization.
The final optimized floorplans in Fig. 14 show that our hybrid approach produces design solutions of similar quality to non-machine learning method while reducing computating timet by $70\times$.
@@ VII Conclusion
This work has presented DeepOHeat-v1 to significantly enhance the original DeepOHeat framework through three key innovations. By replacing MLP trunk networks with Kolmogorov-Arnold Networks, we have enabled adaptive learning of multi-scale thermal features, leading to substantial improvements in prediction accuracy for both surface and volumetric power distributions. The employment of separable training has overcome the computational bottleneck in physics-informed training, making high-resolution thermal analysis feasible on standard GPU hardware. Most importantly, our confidence estimator and hybrid optimization framework transform operator learning from a fast but potentially unreliable predictor into a trustworthy design optimization tool for thermal optimization in 3D-IC design. The selective refinement strategy effectively balances efficiency and reliability, achieving accuracy comparable to an optimization engine using high-fidelityreference mesh-based PDE solvers while obtaining significant speedup. This capability is particularly valuable for practical 3D-IC design, where reliable thermal evaluation is crucial to finding optimal designs in design space exploration and signal/power integrity sign-off.
These advances open new possibilities for efficient thermal optimization in increasingly complex 3D-IC and chiplet design. Several important extensions can build upon this foundation. First, extending to transient thermal analysis by incorporating time as an additional input dimension to the trunk networks. Second, addressing full 3D-IC specific challenges including through-silicon via (TSV) thermal modeling, inter-die thermal coupling, and thermal-aware placement considering packaging constraints. Third, validation on larger chip sizes representative of modern designs. Fourth, and critically, our current test cases demonstrate modest temperature gradients; validation on modern high-performance designs with extreme gradients is essential to fully assess the method’s capabilities. Finally, extending our framework to address more practical thermal-aware floorplan optimization problems, including considerations for power delivery networks, timing constraints, and multi-objective optimization scenarios that balance thermal performance with other critical design metrics.
Figure 15: All 10 test power maps from the original DeepOHeat paper used in Section VI.A evaluation. Top row: power distributions. Rows 2-4: temperature predictions from Celsius 3D (reference), DeepOHeat, and DeepOHeat-v1. Bottom rows: error maps for DeepOHeat and DeepOHeat-v1.